OpenStudy (anonymous):

given that y=Ax^b, find the values of A and b. I have a line lny=-0.5(lnx)+6.8

OpenStudy (kira_yamato):

\[\ln y=-0.5(\ln x)+6.8\] \[\ln y + \ln \sqrt{x} = 6.8\] \[\ln \left[ y \sqrt{x}\right] = 6.8\] \[y \sqrt{x} = e^{6.8}\] \[y = e^{6.8} \times x^{-\frac{1}{2}}\] \[A = e^{6.8}, B=-\frac{1}{2}\]

OpenStudy (anonymous):

thank you!